# The Geometry behind the Two-Dimensional Hard-Disk ‘Glass Transition’

### v1

@@ V Particle Dynamics
Let us now consider the dynamics in the systems. As explained in Section III, I obtained MSDs for a range of $\phi$ given a specific value of $R$. That is, I computed $\langle|\boldsymbol{r}(t)|^{2}\rangle$, where $\boldsymbol{r}$ is an individual particle’s displacement from (reduced) time $t=0$ and the angular brackets indicate ensemble averaging. Figure 3a provides examples of such MSD data for $R^{-1}=1.45$, which is close to the experimental $R$ value of Ref. [1] and also a representative sample for the other $R$ considered in this work. Data for which the standard error exceeds 5% of the mean on the last point is not shown. However, it should be noted that data for values of $\phi>0.744$ was studied and that within the error in this data, I observe a plateau to the MSD, which is indicative of strong caging.
Figure 3: Analyzing the dynamics in the system. All the data in this set of graphs was obtained for $R^{-1}=1.45$. (a) The reduced mean squared displacement (MSD) as a function of the reduced time $t$. From red to blue the value of the area fraction $\phi$ decreases from $0.081$, $0.155$, $0.223$, $0.285$, $0.341$, $0.393$, $0.44$, $0.483$, $0.522$, $0.558$, $0.606$, $0.62$, $0.634$, $0.647$, $0.66$, $0.672$, $0.683$, $0.694$, $0.718$, $0.723$, $0.731$, $0.739$, to $0.744$, respectively. The dots show the average and the error bars the standard error of the mean. The dashed black line shows the linear scaling of the MSD for large times. (b) Logarithmic plot of the reduced diffusion coefficient $D$ as a function of $\phi$ for the data in (a). The error is typically much smaller than the symbol size. The red curve shows a power-law fit to the data. The vertical dashed vertical magenta line indicates the inflection point to the change in diffusion. The inset shows the numerical derivative $\partial D/\partial\phi$ as a function of $\phi$. The solid orange curve shows a quintic polynomial fit to this data and the magenta line is the same as in the main panel.
From the selected curves, I extracted the power-law coefficient $\alpha$ that best described the long-time behavior of the MSD, i.e., $\langle|\boldsymbol{r}|^{2}\rangle\propto t^{\alpha}$. The ESI [59] provides full details on the procedure. Whenever this fitting procedure led to an $\alpha\in(0.97,1]$, I used the fit to also obtain the effective, reduced diffusion coefficient $D$. As expected, $D$ decreased as a function of $\phi$, see Fig. 3b. It was found to exhibit power-law decay close to the point where $D$ became vanishingly small. Beyond the associated $\phi$ value, the MSD also typically started to become significantly subdiffusive (not shown here).
@@
My approach can raise questions regarding the quality of the fit and precision with which I obtained $\phi_{a}$, as well as its sensitivity to the procedure of preparing the systems. To address the first of these concerns, I approached finding $\phi_{a}$ using structural characteristics as well, see Appendix D, wherein I use shell percolation [87]. In brief, there is a small, but measurable feature in the way the shell-percolation point changes with the shell width $\epsilon$ as a function of $\phi$. I found excellent agreement between the $\phi$ for which this feature appears and $\phi_{a}$, in the range where $\phi_{a}$ was observed to be nearly constant. Here, I should emphasize that arrested dynamics, as identified through $\phi_{a}$, is not necessarily indicative of glassy dynamics. It may equally result from the system crystallizing or undergoing another type of phase transition, as will become clear in Section VI.
@@
Turning to the second point, it is not clear that the value of $\phi_{a}$ is insensitive to equilibration and preparation. I address this issue by performing simulations with equilibration lengths ranging from $10^{3}$ and $10^{5}$ reduced time units (post formation by growth) for several values of $\phi$ and $R$, see the ESI [59]. The changes across this time range to the structure were nominal. This provides confidence in my analysis of the dynamics and structure.
@@
Having confirmed that the $\phi_{a}$ feature is robust to changes in equilibration time, I consider a final aspect of $D(\phi)$. Using a simple midpoint scheme, I numerically compute the derivative of $D$ as a function of $\phi$, see the inset to Fig. 3b. This data is fitted using a quintic polynomial (orange curve) and from this the inflection point to $\partial D/\partial\phi$ is obtained. I refer to the associated area fraction as $\phi_{i}$ and have indicated it in the inset using a vertical dashed magenta line. In the main panel (Fig. 3b), $\phi_{i}$ appears to be at the edge of where $D(\phi)$ decays slowly. I have suggestively indicated this by using dashing for the red curve to the left of the inflection point, though it should be noted that lin-log representations can be slightly deceptive in their interpretation. I will discuss the inflection-point data further in Section VII.
@@
I close this section by mentioning that I have also briefly considered the decay present in the self-intermediate scattering function (SISF). It proved challenging to extract $\phi_{a}$ from the SISF data. Referencing the experiments of Lozano et al. [1] this is unsurprising. The relaxation times obtained from their SISF follow the Vogel-Fulcher-Tammann (VTF) law [58, 27] exceptionally well. This can be appreciated from their Fig. S3a and the re-analysis of their data in the ESI [59]. The area fraction for which the divergence occurs extracted from the VFT fit is $\phi\approx 0.845$, which is close to where I place the $\phi_{m}$  [^5] Small departures are expected as the two sizes of colloids used have a little polydispersity and the smaller colloids can be slightly beneath the larger colloids, as their center of mass can be closer to the substrate.. That is, the experimental data did not reveal a direct relation between properties of the SISF and the ‘glass transition’ that was observed. In view of this, I did not pursue further finding $\phi_{a}$ using the SISF, also because there were already two different approaches that gave good agreement for the value.
@@ VI State Diagram
Figure 4: State diagram for bidisperse disks in 2D. The data is represented in a size ratio $R$ to area fraction $\phi$ diagram. Small dots with error bars provide the numerical data and the standard error of the mean, respectively. The solid curves provide guides to the eye for the trends. The black dots and light-black curve provide the random-close packing value, labelled ‘rcp’. The red dots and light-red curve shows those configurations for which the average generalized number of edges $n_{\bar{q}}=5.5$. The red curve is thicker than the other ones, as there is more uncertainty as to what its shape is, see main text. The current representation involves a quartic fit near $R\approx 0.83$. The green dots and light-green curves show those systems, for which $\phi$ the value of $n_{\bar{q}}$ is maximized. The blue dots show the points for which the extrapolated dynamics is fully arrested, $\phi_{a}$, as extracted from the mean-squared displacement, see Section III for the procedure. The horizonal, thick dashed, green line provides the value at which the systems fall out of equilibrium. The six numbered magenta dots correspond to the systems, for which I show a representative snapshot at the bottom of the figure, as labelled. In these snapshots the smaller disks are indicated using a lighter shade of red, which is related to the ratio $R$. The large black dot marked “a” indicates the glass-transition value $\phi=0.776$ for $R\approx 0.7$ as obtained in the experiment of Ref. [1].
@@
The black points provide the $\phi_{m}$ upper bound for having a (disordered) system that is close packed. Note that for sufficiently high $R\approx 1$, the system is ordered and the notion of ‘random’ in rcp is abused, as snapshot 5 clearly reveals crystalline packing. The algorithm of Desmond and Weeks [75] itself does not check for a disordered structure, only for close packing, but I will use it throughout, nonetheless. The light-black curve is a high-order polynomial fit to the data points and serves to guide the eye. The width of this (and other) curve(s) in the diagram corresponds roughly to the maximum deviation from the mean (standard error of the mean).
@@
I correlate the presence of crystallinity throughout the system with $n_{\bar{q}}>5.5$; see Fig. 2d and its inset. For $R\gtrsim 0.83$ there are two intersection points for $n_{\bar{q}}=5.5$, this leads to two distinct branches which must meet up in a single point. It proved difficult to find this point even for my fine-meshed $R$ analysis, I therefore fit the data points closest to $R\approx 0.83$ using a polynomial containing a linear, square, and quartic term. This together with a polynomial fit of the data closer to $R=1$, see the ESI [59], leads to the red envelope presented in Fig. 4. The curve for this envelope is thicker than that used for the other guides to the eye in the diagram, as there is more uncertainty on it, especially close to the ‘critical point’ in terms of $R$. Note that because there is low diffusivity above $\phi_{a}$ the analogy to an equilibrium critical point should not be taken too far.
@@
Snapshots 5 and 6 reveal that the system is crystalline inside of this envelope, while snapshots 3 and 4 reveal disorder. Note that shapshot 3 has many crystalline areas in it, though the overall structure is clearly not that of an ordered solid. Analyzing the sample using $n_{\bar{q}}$ is thus useful to gain insight into bulk features. In Appendix C, I study the pressure as a function of $\phi$. I use this to show that there is a reasonable agreement between the crystalline-phase coexistence density, as measured using the van-der-Waals loop, and the lower $\phi$-branch that results from my (intuited) $n_{\bar{q}}=5.5$ crystallization criterion. There is also satisfactory agreement with $\phi$ coexistence values previously reported in the literature for bidisperse hard-disk systems [89], given that my equilibration procedure is different from that of Huerta et al. This leads me to conclude that crystallization region is accurately located in the diagram.
@@
Figure 4 also shows the value of $\phi$, for which $n_{\bar{q}}$ assumes its maximum (green data). Note that the data clearly has two trends to it, which were individually fitted using a sigmoidal curve. This leads to the solid, light-green curve that also serves to guide the eye. Above $R\approx 0.83$, I find that $n_{\bar{q}}$ shows a descending trend with decreasing $R$. While below this value of $R$, $n_{\bar{q}}$ increases. Transition between these two regimes coincides with the closing of the $n_{\bar{q}}=5.5$ (red) envelope within the error. The two trends are caused by the same effect. Above the $\phi$ value associated with $\max_{\phi}\!\left(n_{\bar{q}}\right)$, the average neighborhood becomes increasingly anisotropic to accommodate the target area fraction. This could be caused by a global distortion of neighborhoods or by the incorporation of defects. I will provide additional evidence in Section VII.
@@
For the systems that crystallize — within the $n_{\bar{q}}=5.5$ envelope — the lower the value of $R$, the more freedom there is to distort neighborhoods, due to the presence of smaller particles. Thus, the value of $\phi$ for which distortion occurs, decreases with decreasing $R$. Outside of the envelope, the system is better characterized by disorder, which is occasionally interspersed with crystalline patches, see also snapshots 1, 2, and 4. Therefore, the smaller the value of $R$, the greater the area fraction must be to begin to see distortions in the neighborhoods of the small particles, which contribute 50% to the value of $n_{\bar{q}}$. I also examined the ‘critical points’ that follow from the intersection of the green curves and from the peak (in $R$) of the red envelope, their location is $\phi\approx 0.777$ $R^{-1}\approx 1.20$ and $\phi\approx 0.785$ $R^{-1}\approx 1.21$, respectively. For $R\lesssim 0.83$, the system loses long-ranged orientational order, see the ESI [59], which leads to the change in trend.
@@
I use the value of $\phi\approx 0.777$ as a tentative indication of where the system appears to fall out of equilibrium, as indicated using the dashed green line that covers the entire $R$ range. I will motivate this extended range more carefully in the upcoming Section VII. Note that this value of $\phi$ is extremely close to what was reported for the experimental system of Lozano et al. [1], which found a ‘glass transition’ at $\phi\approx 0.776$ and $R\approx 0.70$.
@@
Interestingly, when I analyze the value for which the (extrapolated) diffusion coefficient vanishes, $\phi_{a}$, I find that this approaches $\phi=0.777$ for sufficiently low $R$. That is, $\phi_{a}$ matches the constant dashed green curve well for $R\lesssim 0.8$ or equivalently $R^{-1}\gtrsim 1.25$. For $R\gtrsim 0.833$ ($R^{-1}\lesssim 1.2$), I further observe that $\phi_{a}$ accurately tracks the lower part of the red $n_{\bar{q}}=5.5$ envelope. This makes sense, as diffusion in a crystal is mediated by interstitial mechanisms [90], which means that it is generally much slower than diffusion in a fluid. My fitting procedure for $\phi_{a}$ picks up on this difference in dynamics. The more surprising result is the near constancy of $\phi_{a}$ over a large range in $R$, which is what I will turn to next.

### v2

@@ V Particle Dynamics
As explained in Section III, I obtained MSDs for a range of $\phi$ given a specific value of $R$. That is, I computed $\langle|\boldsymbol{r}(t)|^{2}\rangle$, where $\boldsymbol{r}$ is an individual particle’s displacement from (reduced) time $t=0$ and the angular brackets indicate ensemble averaging. Figure 5a provides examples of such MSD data for $R^{-1}=1.45$, which is close to the experimental $R$ value of Ref. [1], as well as a representative sample for the other $R$ considered in this work. Data for which the standard error exceeds 5% of the mean on the last point is not shown. However, it should be noted that data for values of $\phi>0.744$ (not shown here) was studied and that within the error in this data, I observe a plateau to the MSD, which is indicative of strong caging.
Figure 5: Analyzing the dynamics in the system. All the data in this set of graphs was obtained for $R^{-1}=1.45$. (a) The reduced mean squared displacement (MSD) as a function of the reduced time $t$. From red to blue the value of the area fraction $\phi$ increases from $0.081$, $0.155$, $0.223$, $0.285$, $0.341$, $0.393$, $0.44$, $0.483$, $0.522$, $0.558$, $0.606$, $0.62$, $0.634$, $0.647$, $0.66$, $0.672$, $0.683$, $0.694$, $0.718$, $0.723$, $0.731$, $0.739$, to $0.744$, respectively. The dots show the average and the error bars the standard error of the mean. The dashed black line shows the linear scaling of the MSD for large times. (b) Logarithmic plot of the reduced diffusion coefficient $D$ as a function of $\phi$ for the data in (a). The error is typically much smaller than the symbol size. The red curve shows a power-law fit to the data. The vertical dashed vertical magenta line indicates the inflection point to the change in diffusion. The inset shows the numerical derivative $\partial D/\partial\phi$ as a function of $\phi$. The solid orange curve shows a quintic polynomial fit to this data and the magenta line is the same as in the main panel.
From the selected curves, I extracted the power-law coefficient $\alpha$ that best described the long-time behavior of the MSD, i.e., $\langle|\boldsymbol{r}|^{2}\rangle\propto t^{\alpha}$. The ESI [64] provides full details on the procedure. Whenever this fitting procedure led to an $\alpha\in(0.97,1]$, I used the fit to also obtain the effective, reduced diffusion coefficient $D$. As expected, $D$ decreased as a function of $\phi$, see Fig. 5b. It was found to exhibit power-law decay close to the point where $D$ became vanishingly small. Beyond the associated $\phi$ value, the MSD also typically started to become significantly subdiffusive (again, not shown here).
@@
My approach can raise questions regarding the quality of the fit and precision with which I obtained $\phi_{a}$, as well as its sensitivity to the procedure of preparing the systems. To address the first of these concerns, I approached finding $\phi_{a}$ using structural characteristics as well, see Appendix D, wherein I use shell percolation [118]. In brief, there is a small, but measurable feature in the way the shell-percolation point changes with the shell width $\epsilon$ as a function of $\phi$. Note that there is an analogy to this measure and the way contact networks are analyzed in granular packings under compression [82, 3]. Namely, under uniform compression which only slightly perturbs the system, one would expect (on average) contacts to form there where the distances between particles are smallest. When the $\epsilon$-based percolation point changes as a function of $\phi$, this can indicate the presence of nascent load-baring networks, which is what was recently studied using topological analysis in a granular system [3].
@@
I found excellent agreement between the $\phi$ for which this feature appears and $\phi_{a}$, in the range where $\phi_{a}$ was observed to be nearly constant. Here, I should emphasize that arrested dynamics, as identified through $\phi_{a}$, is not necessarily indicative of glassy dynamics. It may equally result from the system crystallizing or undergoing another type of phase transition, e.g., a topological one.
@@
Turning to the second point, it is not clear that the value of $\phi_{a}$ is insensitive to equilibration and preparation. I addressed this issue by performing simulations with equilibration lengths ranging from $10^{3}$ and $10^{5}$ reduced time units (post formation by growth) for several values of $\phi$ and $R$, see the ESI [64]. The changes across this time range to the structure proved to be nominal. This provides confidence in my analysis of both the dynamics and structure around $\phi\approx 0.77$ (and beyond) and contextualizes my earlier statement regarding the quality of the equilibration and sampling.
@@
Having confirmed that the $\phi_{a}$ feature is robust to changes in equilibration time, I consider a final aspect of $D(\phi)$. Using a simple midpoint scheme, I numerically compute the derivative of $D$ as a function of $\phi$, see the inset to Fig. 5b. This data is fitted using a quintic polynomial (orange curve) and from this the inflection point to $\partial D/\partial\phi$ is obtained. I refer to the associated area fraction as $\phi_{i}$ and have indicated it in the inset to Fig. 5b using a vertical dashed magenta line.
In the main panel (Fig. 5b), $\phi_{i}$ appears to be at the edge of where $D(\phi)$ decays slowly. I have suggestively indicated this by using dashing for the red curve to the left of the inflection point, though it should be noted that lin-log representations can be slightly deceptive. The physical interpretation of this point is that $D$ transitions from a change that appears to plateau, i.e., a tendency of $D$ to decrease linearly with $\phi$, to accelerated decrease. This can be a signal of a new pathway to $D$ reduction becoming available in the fluid. I will discuss the inflection-point data further in Section VII.
@@
I close this section by mentioning that I have also briefly considered the decay present in the self-intermediate scattering function (SISF). It proved challenging to extract $\phi_{a}$ from the SISF data. Referencing the experiments of Lozano et al. [1] this is unsurprising. The relaxation times obtained from their SISF follow the Vogel-Fulcher-Tammann (VTF) law [63, 31] exceptionally well. This can be appreciated from their Fig. S3a and the re-analysis of their data in the ESI [64]. The area fraction for which the divergence occurs extracted from the VFT fit is $\phi\approx 0.845$, which is close to where I place the $\phi_{m}$ ($\phi$ for RCP)  [^10] Small departures are expected as the two sizes of colloids used have a little polydispersity and the smaller colloids can be slightly beneath the larger colloids, as their center of mass can be closer to the substrate.. That is, the experimental data did not reveal a direct relation between properties of the SISF and the glass transition that was experimentally observed. In view of this, I did not pursue further finding $\phi_{a}$ using the SISF, also because there were already two different approaches that were internally consistent.
@@ VI State Diagram
Figure 6: State diagram for bidisperse disks in 2D. The data is represented in a size ratio $R$ to area fraction $\phi$ diagram. Small dots with error bars provide the numerical data and the standard error of the mean, respectively. The solid curves provide guides to the eye for the trends. The black dots (error smaller than symbol size) and light-black curve provide the random-close packing value, labelled “RCP”. The red dots and light-red curve shows those configurations for which the average generalized number of edges $n_{\bar{q}}=5.5$. The red curve is thicker than the other ones, as there is more uncertainty as to what its shape is, see main text. The current representation involves a quartic fit near $R\approx 0.83$, see the ESI [64]. The green dots and light-green curves show those systems, for which $\phi$ the value of $n_{\bar{q}}$ is maximized; sigmoidal fits were used here, see the ESI [64]. The blue dots show the points for which the extrapolated dynamics is fully arrested, $\phi_{a}$, as extracted from the MSD. These points partially overlap with the red dots along the bottom red curve. The horizonal, thick-dashed orange line provides the value at which the systems fall out of equilibrium for $R\lesssim 0.8$ and is extrapolated to higher values of $R$. The six numbered magenta dots correspond to the systems, for which representative snapshots are shown at the bottom of the figure, as labelled. In these snapshots the larger disks are indicated in red and the smaller disks ones using blue. The left-bottom corner for each snapshot shows a part of the 2D scattering data. This is provided on a logarithmic color scale, see the ESI [64] for details. The large black dot marked “a” indicates the glass-transition value $\phi=0.776$ for $R\approx 0.698$ as obtained in the colloidal experiment of Ref. [1]. The black bar marked “b” indicates the range of values for the jamming point that was obtained in the granular experiments of Refs. [2, 3], $\phi\approx 0.773$ to $0.777$, with $R\approx 0.714$.
@@
The black points provide the $\phi_{m}$ upper bound for having a (disordered) system that is close packed. Note that for sufficiently high $R\approx 1$, the system is ordered and the notion of ‘random’ in RCP is abused, as snapshot 5 clearly reveals crystalline packing. The algorithm of Desmond and Weeks [109] does not check for disorder, only for close packing, but I will use the term RCP throughout, nonetheless. The light-black curve is a high-order polynomial fit to the data points and serves to guide the eye. The width of this (and other) curve(s) in the diagram corresponds roughly to the maximum deviation from the mean (standard error of the mean) in the data.
I correlate the presence of crystallinity throughout the system with $n_{\bar{q}}>5.5$. For $R\gtrsim 0.83$ there are two intersection points for $n_{\bar{q}}=5.5$, see Fig. 4d and its inset, which leads to two distinct branches (red points) that must meet up in a single ‘critical’ point. In Appendix C, I show that $n_{\bar{q}}\geq 5.5$ is where the bond-orientational order parameter predicts a hexagonal crystal. Snapshots 5 and 6 also reveal that the system is crystalline inside of this envelope, while snapshots 3 and 4 reveal disorder. Note that shapshot 3 has several crystalline areas in it, though the overall structure is clearly not that of an ordered solid, as also evidenced by the scattering profile provided in the insets to the snapshots. Analyzing the sample using $n_{\bar{q}}$ is thus meaningful.
@@
It proved difficult to find the ‘critical’ point even for my fine-meshed $R$ analysis; I therefore fitted the data points closest to $R\approx 0.83$ using a polynomial of the form $R=A\phi+B\phi^{2}+C\phi^{4}$ with $A$, $B$, and $C$ fit coefficients. This together with another quartic polynomial fit of $\phi$ in terms of $R$ on the data closer to $R=1$, see the ESI [64], leads to the red curve (envelope to the crystalline region) presented in Fig. 6. The curve for this envelope is thicker than that used for the other guides to the eye in the diagram, as there is more uncertainty on it, especially close to the ‘critical point’ in terms of $R$. Here, I use this wording, because the value of $R\approx 0.83$ appears to delimit the crystalline region in this manner. Note that because there is low diffusivity close to $\phi_{a}$, the analogy to an equilibrium critical point of a temperature-density phase diagram is tenuous. Nonetheless, I will argue for a topological transition se tting in from this point.
@@
Appendix C also shows results for the pressure as a function of $\phi$. I use this to demonstrate that there is a reasonable agreement between the crystalline-phase coexistence density, as measured using the van-der-Waals loop, and the lower $\phi$-branch resulting from the $n_{\bar{q}}~=~5.5$ criterion. There is also satisfactory agreement with $\phi$ coexistence values previously reported in the literature for bidisperse hard-disk systems [120]. All of this together leads me to conclude that crystallization region is accurately located in the diagram. Finally, obtaining the hexatic transition was not a goal in this study and considering the system sizes required to accurately determine this value [121, 122] is not valuable or needed.
@@
Figure 6 also shows the value of $\phi$, for which $n_{\bar{q}}$ assumes its maximum (green data). Note that the data clearly has two trends to it, which were individually fitted using a sigmoidal curve. This leads to the solid, light-green curve that also serves to guide the eye. Above $R\approx 0.83$, I find that $n_{\bar{q}}$ shows a descending trend with decreasing $R$, while below this value of $R$, $n_{\bar{q}}$ increases with decreasing $R$. The transition between these two regimes coincides with the closing of the $n_{\bar{q}}=5.5$ (red) envelope within the error. The two trends are caused by the same effect. Above the $\phi$ value associated with $\max_{\phi}\!\left(n_{\bar{q}}\right)$, the average neighborhood becomes increasingly anisotropic to accommodate the target area fraction. This could be caused by a global distortion of neighborhoods or by the incorporation of defects. I will provide additional evidence in Section VII.
@@
The trends displayed by the green points may be explained as follows. For the systems that crystallize — within the $n_{\bar{q}}=5.5$ envelope — the lower the value of $R$, the more freedom there is to distort neighborhoods, due to the presence of smaller particles. Thus, the value of $\phi$ for which distortion occurs, decreases with decreasing $R$. Outside of the envelope, i.e., for $R\lesssim 0.83$, the system is better characterized by disorder, which is occasionally interspersed with crystalline patches, see also snapshots 1, 2, and 4. Therefore, the smaller the value of $R$, the greater the area fraction must be to begin to see distortions in the neighborhoods of the small particles, which contribute 50% to the value of $n_{\bar{q}}$.
@@
Lastly, I examined the point that follows from the intersection of the green curves, which is located at $\phi\approx 0.777$ with $R^{-1}\approx 1.20$. The area fraction belonging to this intersection value is indicated over the entire $R$ range using the horizonal, thick-dashed orange line. I will motivate this extended range more carefully in the upcoming Section VII. This value of $\phi$ is extremely close to what was reported for the experimental systems of Lozano et al. [1] and Refs. [2, 3], see point “a” and line “b” in the state diagram 6, respectively.
@@
Interestingly, the values of $\phi$ for which the (extrapolated) diffusion coefficient vanishes, $\phi_{a}$, correspond well to $\phi=0.777$ for sufficiently low $R$. That is, $\phi_{a}$ matches the constant dashed orange curve well for $R\lesssim 0.8$ or equivalently $R^{-1}\gtrsim 1.25$. For $R\gtrsim 0.833$ ($R^{-1}\lesssim 1.2$), I further observe that $\phi_{a}$ accurately tracks the lower part of the red $n_{\bar{q}}=5.5$ envelope. This makes sense, as diffusion in a crystal is mediated by interstitial mechanisms [123], which means that it is generally much slower than diffusion in a fluid. My fitting procedure for $\phi_{a}$ picks up on this difference in dynamics. The more surprising result is the near constancy of $\phi_{a}$ over a large range in $R$, which is what I will turn to next.